23. On January 1, 2018, Legion Company sold $200,000 of 10% ten-year bonds. Interest is payable semiannually on June 30 and December 31 . The bonds were sold for S 177,000, priced to yield 12% Legion records interest at the effective rate. Legion should report bond interest expense for the six months ended June 30, 2018, in the amount of: A) $10,000. B) $10,620. C) $12,000. D) $8,850. Problems 24-26 are a part of a multi-step problem using the following information Lopez Plastics Co. (LPC) issued callable bonds on January 1, 2018. LPCs accountant has projected the following amortization schedule from issuance until maturity Cash interest Effective Decrease in Outstanding Date 1/1/2018 6/30/2018 12/31/2018 6/30/2019 12/31/2019 S 7,000 7,000 7,000 7,000 7,000 7,000 7,000 7,000 s 207,020 206.230 205,417 204,580 203,717 202,829 201,913 200,971 200,000 s 789 813 837 863 S 6,211 6,187 6,163 6,137 6,112 6,085 6,057 6,029 12/31/2020 6/30/2021 12/31/2021 915 943 971 What is the annual stated interest rate on the bonds? a) 7% B) 3.5% D) None of the answer choices is correct 25. What is the annual effective interest rate on the bonds? A) 6% B) 3.5% C) 796 D) 3% 26. LPC calls the bonds at 103 immediately after the interest payment on 12/31/2019 and retires them. What gain or loss, ifany, would LPC record on this date? A) $3,717 gain B) S6,000 loss C No gain or loss D)$2,283 loss 6
Answer #1

Solution 23:

Face value of bond = $200,000

Issue price of bond = $177,000

Coupon rate = 10%

Effective interest rate = 12%

Bond interest expense to be reported for 6 months ended June 30, 2018 = $177,000*12%*6/12 = $10,620

Hence option B is correct.

Solution 24:

Face value of bond = Maturity value of bond = $200,000

Semi annual interest payment = $7,000

Semi annual interest rate = $7,000/$200,000 = 3.50%

Annual stated interest rate on bond = 3.50*2 = 7%

Hence option A is correct.

Solution 25:

Issue price of bond = $207,020

Effective interest for first 6 months = 6,211

Semi annual effective interest rate = $6,211 / $207,020 = 3%

Effective annual interest rate = 3*2 = 6%

Hence option A is correct.

Solution 26:

Balance of bond on 31.12.2019 = $203,717

Reitrement value of bond = $200,000*103% = $206,000

Loss on retirement = $206,000 - $203,717 = $2,283

Hence Option D is correct.